If tan(theta/2) = 2/3 then sec theta = ?
a).13/5 b). 3/13 c). 5/13 d). 12/13

Grade:11

2 Answers

Suraj Kumar
6 years ago
Using formula Tan^(theta/2)=(1-cos□)/(1+cos□)where ^=square and □=theta4/9=(1-cos□)/(1-cos□)By cross multiplication We get 13cos□=5Cos□=5/13Sec□=13/5So (a)
Arjun Raj
4 years ago
tan(x/2)=2/3.                                                              cosx ={1-tan2(x/2)}/{1+tan2(x/2)} .                              cosx={1-(4/9)}/{1+(4÷9)}.                                               cosx=5/13.           ............                                         secx=13/5